OpenStudy (anonymous):

(x+4/3)=2..... x+4 is over 3

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x+4}{3}=2\] \[\Large x+4=2*3\] \[\Large x+4=6\] \[\Large x=6-4\] \[\Large x=2\]